# Heapr: Hessian-based Efficient atomic expert Pruning in Output Space

### v1

@@ Abstract
Mixture-of-Experts (MoE) architectures in large language models (LLMs) deliver exceptional performance and reduced inference costs compared to dense LLMs. However, their large parameter counts result in prohibitive memory requirements, limiting practical deployment. While existing pruning methods primarily focus on expert-level pruning, this coarse granularity often leads to substantial accuracy degradation. In this work, we introduce Heapr, a novel pruning algorithm that decomposes experts into smaller, indivisible atomic experts, enabling more precise and flexible atomic expert pruning. To measure the importance of each atomic expert, we leverage second-order information based on principles similar to Optimal Brain Surgeon (OBS) theory. To address the computational and storage challenges posed by second-order information, Heapr exploits the inherent properties of atomic experts to transform the second-order information from expert parameters into that of atomic expert parameters, and further simplifies it to the second-order information of atomic expert outputs. This approach reduces the space complexity from $O(d^{4})$, where $d$ is the model’s dimensionality, to $O(d^{2})$. Heapr requires only two forward passes and one backward pass on a small calibration set to compute the importance of atomic experts. Extensive experiments on MoE models, including DeepSeek MoE and Qwen MoE family, demonstrate that Heapr outperforms existing expert-level pruning methods across a wide range of compression ratios and benchmarks. Specifically, Heapr achieves nearly lossless compression at compression ratios of $20\%\sim 25\%$ in most models, while also reducing FLOPs nearly by $20\%$. The code can be found at https://github.com/LLIKKE/HEAPr.
@@ 1 Introduction
Mixture-of-experts (MoE) models have recently emerged as a promising alternative to dense large language models (LLMs), replacing dense feed-forward layers with sparsely activated experts and dynamic routing. This design allows MoE models to match or surpass the performance of dense LLMs while activating only a fraction of parameters during inference (Fedus et al., 2022; Zhu et al., 2024; Liu et al., 2024), making them particularly attractive for large-scale, concurrent deployment. However, while sparse activation reduces computational cost, it exacerbates memory requirements. For example, DeepSeek-V3 (Liu et al., 2024) activates only 37B parameters per inference, yet all 671B parameters must still be stored in GPU memory, resulting in prohibitively high deployment costs. Notably, MoE layers typically account for over $97\%$ of total model parameters, they represent the dominant storage bottleneck. Therefore, compressing MoE layers becomes critical to overcoming inference inefficiency and making deployment feasible in resource-constrained environments.
@@
Model pruning has been widely explored as an effective compression strategy to reduce storage and improve efficiency. Yet a fundamental trade-off persists: fine-grained pruning typically preserves accuracy but yields limited speedups on hardware, whereas coarse-grained pruning directly enables acceleration but often incurs obvious accuracy loss. Within MoE models, parameter sparsification (Xie et al., 2024) faces similar limitations, as hardware inefficiencies constrain its practical
@@
Figure 1: Illustration of atomic expert-level pruning, which removing the $t$-th column from the $W_{gate}$ and $W_{up}$ matrices, and the corresponding $t$-th row from the $W_{down}$ matrix.
@@
benefits. Consequently, recent research has shifted toward expert-level pruning, offering more direct gains in both acceleration and memory reduction. Existing expert-level approaches at this level can be broadly divided into expert dropping and expert merging. Expert dropping methods (Lu et al., 2024; Huang et al., 2025) completely remove experts deemed unimportant, but relying solely on calibration to discard entire experts risks losing valuable complementary expertise, consequently often leading to notable performance degradation. Expert merging methods (Li et al., 2024; Chen et al., 2025; Huang et al., 2025) instead aim to consolidate functionally similar experts to more effectively preserve overall model capacity. However, their clustering-based similarity measures are notoriously unstable, and naive merging strategies (e.g., averaging or frequency-based weighting) often introduce destructive parameter conflicts, resulting in suboptimal and inefficient outcomes. To alleviate these critical conflicts, recent decomposition-based approaches (Li et al., 2025b; Gu et al., 2025) represent individual experts as a mixture of shared and specialized components. While this advanced framework helps to preserve model capacity, it still requires computationally expensive decomposition and merging operations, and unfortunately still incurs a non-negligible accuracy loss.
@@
  • We conduct extensive experiments on DeepSeekMoE-16B-Base, Qwen1.5-MoE-A2.7B-Chat, Qwen2-57B-A14B, and Qwen3-30B-A3B across diverse benchmarks. Heapr outperforms current SOTA methods and achieves nearly lossless compression at compression ratios of $20\%–25\%$ in most models, while also reducing FLOPs by nearly $20\%$.
@@
Model compression for MoE architectures has recently attracted growing attention due to the remarkable performance of MoE models. MoE-Pruner (Xie et al., 2024) performs weights sparsification based on activation magnitude, weight magnitude, and router importance, yet its acceleration is hardware-dependent and relies on distillation to recover accuracy. Expert-level pruning has been more extensively explored due to its hardware-friendly acceleration. NAEE (Lu et al., 2024) selects a subset of experts to minimize calibration error, but this can lead to overfitting and the loss of specialized knowledge. Similarly, MoE-I2 (Yang et al., 2024) combines expert pruning with low-rank decomposition, yet requires additional fine-tuning for recovery. To alleviate such issues, expert merging methods aim to retain similar experts rather than discarding them. MC-SMoE (Li et al., 2024) merges experts by clustering based on routing policies, and HC-MoE (Chen et al., 2025) does so by grouping experts with similar outputs. However, limited expert similarity makes merging prone to parameter conflicts. To further exploit redundancy, $D^{2}$-MoE (Gu et al., 2025) constructs a shared expert via weighted combinations and compresses residuals through low-rank decomposition, while Sub-MoE (Li et al., 2025a) applies SVD to extract a shared subspace across experts, both of which require computationally expensive decomposition and merging operations. We decompose the expert into atomic experts and propose Heapr, a method that measures atomic expert importance by utilizing a second-order approximation to assess the importance of atomic experts. This approach enables more flexible pruning units and provides a highly efficient algorithmic, preserving model performance while eliminating the need for retraining.
@@
Directly computing the full Hessian in deep neural networks is practically infeasible. Existing OBS methods to adopt significant approximations. For instance, K-FAC approximation (Martens & Grosse, 2015) provides an efficient approximation of second-order information and Hessians are computed layer-wise to guide pruning (Dong et al., 2017; Frantar & Alistarh, 2022; Frantar et al., 2023). Previous work (Singh & Alistarh, 2020) shows that the Fisher information matrix serves as a reliable Hessian estimate and allows for more efficient computation. Some apply OBS to structured pruning (Yu et al., 2022), but these efforts are limited as they consider only the trace of the Hessian.
@@ 3 Method
The MoE architecture has been widely adopted in LLMs as a replacement for the dense feed-forward network layer, which effectively increases the model capacity while reducing the number of activated parameters. Formally, given an input token representation $\mathbf{x}\in\mathbb{R}^{d_{\text{model}}}$, the MoE output is defined as:
@@
where $\bm{W}^{\mathrm{gate}}\in\mathbb{R}^{E\times d_{\text{model}}}$ produces router scores and $\operatorname{Top}\text{-}\kappa(\cdot)$ denotes the router function that selects the top-$\kappa$ experts. Each expert $E_{i}(\cdot)$ is a gated feed-forward block:
@@ 3.2 Atomic expert Importance Analysis in the Output Space
$\Delta\ell=\frac{1}{2}\delta\Theta^{T}\bm{H}\delta\Theta=\frac{1}{2}\sum_{i=1}^{d_{\text{inter}}}(\delta\Theta^{(i)})^{T}\bm{H}^{(i)}\delta\Theta^{(i)}$ (8)
here, $\Theta\in\mathbb{R}^{3d_{\text{model}}\cdot d_{\text{inter}}}$ denotes the parameters of a given expert, while $\bm{H}$ is the corresponding Hessian matrix, which has a space complexity of $\mathcal{O}\left((3d_{\text{model}}\cdot d_{\text{inter}})^{2}\right)$. The matrix $\bm{H}^{(i)}$ represents the Hessian for the $i$-th atomic expert. This decomposition leads to a significant reduction in the complexity of summing over the Hessians $\sum_{i=1}^{d_{\text{inter}}}\bm{H}^{(i)}$, which is reduced to $\mathcal{O}\left((3d_{\text{model}})^{2}\cdot d_{\text{inter}}\right)$. However, the resulting Hessian matrix computation remains unacceptable due to its high computational and storage cost. To further alleviate the bottleneck, we introduce a second optimization that shifts pruning constraints from the parameter space of atomic experts to their output space, where the constraint in equation 2 can be equivalently enforced in the atomic expert output domain. Specifically, pruning the atomic expert $\mathbf{e}_{\mathcal{P}}$ is equivalent to imposing: $\mathbf{e}_{\mathcal{P}}(\Theta_{\mathcal{P}}+\delta\Theta_{\mathcal{P}})=\mathbf{0},$ where $\Theta_{\mathcal{P}}\in\mathbb{R}^{3d_{\text{model}}}$ denotes the parameters of the pruned atomic expert. Since the atomic expert functions are not optimized with respect to the parameters $\Theta_{\mathcal{P}}$ through gradient descent, applying a Taylor expansion of the atomic expert functions around $\Theta_{\mathcal{P}}$ results in the first-order term dominating, yielding:
$\mathbf{e}_{\mathcal{P}}(\Theta_{\mathcal{P}}+\delta\Theta_{\mathcal{P}})\approx\mathbf{e}_{\mathcal{P}}(\Theta_{\mathcal{P}})+\bm{J}_{\mathcal{P}}\delta\Theta_{\mathcal{P}}=\mathbf{0},$ (9)
where $\bm{J}_{\mathcal{P}}\in\mathbb{R}^{d_{\text{model}}\times 3d_{\text{model}}}$ denotes the Jacobian of $\mathbf{e}_{\mathcal{P}}(\Theta_{\mathcal{P}})$. This leads to the following problem:
$\min_{\Theta_{\mathcal{P}}}\frac{1}{2}\sum_{i=1}^{d_{\text{inter}}}(\delta\Theta^{(i)})^{T}\bm{H}^{(i)}\delta\Theta^{(i)}\quad\text{s.t.}\quad\bm{J}_{\mathcal{P}}\,\delta\Theta_{\mathcal{P}}+\mathbf{e}_{\mathcal{P}}=\mathbf{0}.$ (10)
To solve the problem in equation 10 , we consider the LLMs trained with a negative log-likelihood loss $\ell$ (e.g., cross-entropy loss). In this setting, the Fisher Information Matrix $\bm{F}$ is equivalent to the expected Hessian (Bishop & Nasrabadi, 2006), providing a computationally efficient alternative:
@@
where $\ell$ is the sample-wise loss. Previous work (Singh & Alistarh, 2020) has shown that for well-converged neural networks, a few hundred representative samples are already sufficiently reliable to estimate $\mathbb{E}\left[{\bm{H}}\right]$. Expanding the gradient of $\ell$ with respect to the parameters gives $\nabla_{\Theta_{\mathcal{P}}}\ell=\bm{J}_{\mathcal{P}}^{\top}\mathbf{g}_{\mathcal{P}},$ where $\mathbf{g}_{\mathcal{P}}\in\mathbb{R}^{d{\text{model}}}$ is the gradient of the loss with respect to the pruned atomic expert output $\mathbf{e}_{\mathcal{P}}$. Substituting this expression into the objective equation 10 yields the expected loss increase when pruning the atomic expert $\mathbf{e}_{\mathcal{P}}$, with $\delta\Theta^{(i)}=\mathbf{0}$ for all atomic experts not pruned:
$\Delta\ell=\tfrac{1}{2}\,\delta\Theta_{\mathcal{P}}^{T}\mathbb{E}\left[{\bm{H}}_{\mathcal{P}}\right]\,\delta\Theta_{\mathcal{P}}=\tfrac{1}{2}\,\mathbf{e}_{\mathcal{P}}^{\top}\,\mathbb{E}[\mathbf{g}_{\mathcal{P}}\mathbf{g}_{\mathcal{P}}^{\top}]\,\mathbf{e}_{\mathcal{P}}.$ (12)
@@
$s=\tfrac{1}{2}\,\mathbf{e}_{\mathcal{P}}^{\top}\,\mathbb{E}[\mathbf{g}_{\mathcal{P}}\mathbf{g}_{\mathcal{P}}^{\top}]\,\mathbf{e}_{\mathcal{P}},$ (13)

### v2

# arXiv:2509.22299v2[cs.LG] 28 Feb 2026

@@ Abstract
Mixture-of-Experts (MoE) architectures in large language models (LLMs) deliver exceptional performance and reduced inference costs compared to dense LLMs. However, their large parameter counts result in prohibitive memory requirements, limiting practical deployment. While existing pruning methods primarily focus on expert-level pruning, this coarse granularity often leads to substantial accuracy degradation. In this work, we introduce Heapr, a novel pruning algorithm that decomposes experts into smaller, indivisible atomic experts, enabling more precise and flexible atomic expert pruning. To measure the importance of each atomic expert, we leverage second-order information based on principles similar to the Optimal Brain Surgeon theory. To address the computational and storage challenges posed by second-order information, Heapr exploits the inherent properties of atomic experts to transform the second-order information from expert parameters into that of atomic expert parameters, and further simplifies it to the second-order information of atomic expert outputs. This approach reduces the space complexity from $\mathcal{O}(d^{4})$, where $d$ is the model’s dimensionality, to $\mathcal{O}(d^{2})$. Heapr requires only two forward passes and one backward pass on a small calibration set to compute the importance of atomic experts. Extensive experiments on MoE models, including DeepSeek MoE and Qwen MoE family, demonstrate that Heapr outperforms existing expert-level pruning methods across a wide range of pruning ratios and benchmarks. Specifically, Heapr achieves nearly lossless compression at pruning ratios of $20\%\sim 25\%$ in most models, while also reducing FLOPs by nearly $20\%$. The code can be found at https://github.com/LLIKKE/HEAPr.
@@ 1 Introduction
Mixture-of-Experts (MoE) models have recently emerged as a promising alternative to dense large language models (LLMs), replacing dense feed-forward layers with sparsely activated experts and dynamic routing. This design allows MoE models to match or surpass the performance of dense LLMs while activating only a fraction of parameters during inference (Fedus et al., 2022; Zhu et al., 2024; Liu et al., 2024a), making them particularly attractive for large-scale, concurrent deployment. However, while sparse activation reduces computational cost, it exacerbates memory requirements. For example, DeepSeek-V3 (Liu et al., 2024a) activates only 37B parameters per inference, yet all 671B parameters must still be stored in GPU memory, resulting in prohibitively high deployment costs. Notably, MoE layers typically account for over $97\%$ of total model parameters, and they represent the dominant storage bottleneck. Therefore, compressing MoE layers becomes critical to overcoming inference inefficiency and making deployment feasible in resource-constrained devices.
@@
Model pruning has been widely explored as an effective compression strategy to reduce storage and improve efficiency. Yet a fundamental trade-off persists: fine-grained pruning typically preserves accuracy but yields limited speedups on hardware, whereas coarse-grained pruning directly enables acceleration but often incurs obvious accuracy loss. Within MoE models, parameter sparsification (Xie et al., 2024) faces similar limitations, as hardware inefficiencies constrain its practical benefits.
@@
Figure 1: Illustration of atomic expert-level pruning, which removes the $t$-th column from the $W_{gate}$ and $W_{up}$ matrices, and the corresponding $t$-th row from the $W_{down}$ matrix.
@@
Consequently, recent research has shifted toward expert-level pruning, offering more direct gains in both acceleration and memory reduction. Existing expert-level approaches at this level can be broadly divided into expert dropping and expert merging. Expert dropping methods (Lu et al., 2024; Huang et al., 2025) completely remove experts deemed unimportant, but relying solely on calibration to discard entire experts risks losing valuable complementary expertise, consequently often leading to notable performance degradation. Expert merging methods (Li et al., 2024; Chen et al., 2025; Huang et al., 2025) instead aim to consolidate functionally similar experts to more effectively preserve overall model capacity. However, their clustering-based similarity measures are notoriously unstable, and naive merging strategies (e.g., averaging or frequency-based weighting) often introduce destructive parameter conflicts, resulting in suboptimal and inefficient outcomes. To alleviate these critical conflicts, recent decomposition-based approaches (Li et al., 2025c; Gu et al., 2025) represent individual experts as a mixture of shared and specialized components. While this advanced framework helps to preserve model capacity, it still requires computationally expensive decomposition and merging operations, and unfortunately still incurs a non-negligible accuracy loss.
@@
  • We conduct extensive experiments on DeepSeekMoE-16B-Base, Qwen1.5-MoE-A2.7B-Chat, Qwen2-57B-A14B, and Qwen3-30B-A3B across diverse benchmarks. Heapr outperforms current SOTA methods and achieves nearly lossless compression at pruning ratios of $20\%–25\%$ in most models, while also reducing FLOPs by nearly $20\%$.
@@ 2 Related Works and Preliminary
Model compression for MoE architectures has recently attracted growing attention due to the remarkable performance of MoE models. MoE-Pruner (Xie et al., 2024) performs weights sparsification based on activation magnitude, weight magnitude, and router importance, yet its acceleration is hardware-dependent and relies on distillation to recover accuracy. Expert-level pruning has been more extensively explored due to its hardware-friendly acceleration. NAEE (Lu et al., 2024) selects a subset of experts to minimize calibration error, but this can lead to overfitting and the loss of specialized knowledge. Similarly, MoE-I2 (Yang et al., 2024) combines expert pruning with low-rank decomposition, yet requires additional fine-tuning for recovery. To alleviate such issues, expert merging methods aim to retain similar experts rather than discarding them. MC-SMoE (Li et al., 2024) merges experts by clustering based on routing policies, and HC-MoE (Chen et al., 2025) does so by grouping experts with similar outputs. However, limited expert similarity makes merging prone to parameter conflicts. EEP (Liu et al., 2024b) uses gradient-free evolutionary search to combine expert dropping and expert merging, cutting SMoE experts and active experts while maintaining or improving downstream performance. To further exploit redundancy, $D^{2}$-MoE (Gu et al., 2025) constructs a shared expert via weighted combinations and compresses residuals through low-rank decomposition, while Sub-MoE (Li et al., 2025a) applies SVD to extract a shared subspace across experts, both of which require computationally expensive decomposition and merging operations. We decompose the expert into atomic experts and propose Heapr, a method that measures importance by utilizing a second-order approximation to assess the importance of atomic experts. This approach enables more flexible pruning units and provides an efficient highly algorithm, preserving model performance while eliminating the need for retraining.
@@
Directly computing the full Hessian in deep neural networks is practically infeasible. Existing OBS methods to adopt significant approximations. For instance, K-FAC approximation (Martens and Grosse, 2015) provides an efficient approximation of second-order information and Hessians are computed layer-wise to guide pruning (Dong et al., 2017; Frantar and Alistarh, 2022; Frantar et al., 2023). Previous work (Singh and Alistarh, 2020) shows that the Fisher information matrix serves as a reliable Hessian estimate and allows for more efficient computation. Some apply OBS to structured pruning (Yu et al., 2022), but these efforts are limited as they consider only the trace of the Hessian.
@@ 3 Method
The MoE architecture has been widely adopted in LLMs as a replacement for the dense feed-forward network layer, which effectively increases the model capacity while reducing the number of activated parameters. Formally, given an input token representation $\mathbf{x}\in\mathbb{R}^{d_{\text{model}}}$, the output of the MoE layer with $N_{\text{exp}}$ experts is defined as:
@@
where $\bm{W}^{\mathrm{gate}}\in\mathbb{R}^{N_{\text{exp}}\times d_{\text{model}}}$ produces router scores and $\operatorname{Top}\text{-}\kappa(\cdot)$ denotes the router function that selects the top-$\kappa$ experts. Each expert $E_{i}(\cdot)$ is a gated feed-forward block:
@@ 3.2 Atomic expert Importance Analysis in the Output Space
$\Delta\ell\approx\frac{1}{2}\delta\Theta^{T}\bm{H}\delta\Theta=\frac{1}{2}\sum_{i=1}^{d_{\text{inter}}}(\delta\Theta^{(i)})^{T}\bm{H}^{(i)}\delta\Theta^{(i)}$ (8)
here, $\Theta\in\mathbb{R}^{3d_{\text{model}}\cdot d_{\text{inter}}}$ denotes the parameters of a given expert, and $\bm{H}$ is the corresponding Hessian matrix with space complexity $\mathcal{O}\big((3d_{\text{model}}\cdot d_{\text{inter}})^{2}\big)$. And each $\bm{H}^{(i)}$ represents the Hessian for the $i$-th atomic expert. This decomposition leads to a significant reduction in the complexity of summing over the Hessians $\sum_{i=1}^{d_{\text{inter}}}\bm{H}^{(i)}$, which is reduced to $\mathcal{O}\left((3d_{\text{model}})^{2}\cdot d_{\text{inter}}\right)$.
However, the resulting Hessian matrix computation remains unacceptable due to its high computational and storage cost. To further alleviate the bottleneck, we introduce a second optimization that reformulates the pruning constraint. The original parameter-space constraint (equation 2) implies that the atomic expert’s output $\mathbf{e}_{\mathcal{P}}(\mathbf{x};\Theta_{\mathcal{P}}+\delta\Theta_{\mathcal{P}})$, where $\Theta_{\mathcal{P}}\in\mathbb{R}^{3d_{\text{model}}}$ denotes the parameters of the atomic expert to be pruned, would be zero for every possible input $\mathbf{x}$. Although theoretically sound, enforcing such a universal constraint is computationally infeasible. This motivates a more targeted reformulation: for a specific token $\mathbf{x}$, what is the minimum loss increase $\Delta\ell(\mathbf{x})$ required to force the expert’s output to zero? To make this question concrete, we impose the per-token constraint $\mathbf{e}_{\mathcal{P}}(\mathbf{x};\Theta_{\mathcal{P}}+\delta\Theta_{\mathcal{P}})=0,$ treating $\mathbf{x}$ as given. Since the atomic expert functions are not optimized with respect to the parameters $\Theta_{\mathcal{P}}$ through gradient descent, applying a Taylor expansion of the atomic expert functions around $\Theta_{\mathcal{P}}$ results in the first-order term dominating, yielding:
$\mathbf{e}_{\mathcal{P}}(\mathbf{x};\Theta_{\mathcal{P}}+\delta\Theta_{\mathcal{P}})\approx\mathbf{e}_{\mathcal{P}}(\mathbf{x};\Theta_{\mathcal{P}})+\bm{J}_{\mathcal{P}}\delta\Theta_{\mathcal{P}}=\mathbf{0},$ (9)
where $\bm{J}_{\mathcal{P}}\in\mathbb{R}^{d_{\text{model}}\times 3d_{\text{model}}}$ denotes the Jacobian of $\mathbf{e}_{\mathcal{P}}(\mathbf{x};\Theta_{\mathcal{P}})$. This leads to the following problem:
$\min_{\delta\Theta_{\mathcal{P}}}\frac{1}{2}\sum_{i=1}^{d_{\text{inter}}}(\delta\Theta^{(i)})^{T}\bm{H}^{(i)}\delta\Theta^{(i)}\quad\text{s.t.}\quad\bm{J}_{\mathcal{P}}\,\delta\Theta_{\mathcal{P}}+\mathbf{e}_{\mathcal{P}}=\mathbf{0}.$ (10)
To solve the problem in equation 10 , we consider the LLMs trained with a negative log-likelihood loss $\ell$ (e.g., cross-entropy loss). In this setting, the Fisher Information Matrix $\bm{F}$ is equivalent to the expected Hessian (Bishop and Nasrabadi, 2006), providing a computationally efficient alternative:
@@
where $\ell$ is the sample-wise loss. Previous work (Singh and Alistarh, 2020) has shown that for well-converged neural networks, a few hundred representative samples are already sufficiently reliable to estimate $\mathbb{E}\left[{\bm{H}}\right]$. Expanding the gradient of $\ell$ with respect to the parameters gives $\nabla_{\Theta_{\mathcal{P}}}\ell=\bm{J}_{\mathcal{P}}^{\top}\mathbf{g}_{\mathcal{P}},$ where $\mathbf{g}_{\mathcal{P}}\in\mathbb{R}^{d{\text{model}}}$ is the gradient of the loss with respect to the pruned atomic expert output $\mathbf{e}_{\mathcal{P}}$. Substituting this expression into the objective equation 10 yields the expected loss increase when pruning the atomic expert $\mathbf{e}_{\mathcal{P}}$, with $\delta\Theta^{(i)}=\mathbf{0}$ for all atomic experts not pruned:
$\tfrac{1}{2}\,\delta\Theta_{\mathcal{P}}^{T}\mathbb{E}\left[{\bm{H}}_{\mathcal{P}}\right]\,\delta\Theta_{\mathcal{P}}\approx\mathbb{E}_{\mathbf{x}\sim D}\left[\tfrac{1}{2}\,\mathbf{e}_{\mathcal{P}}^{\top}\,\mathbb{E}[\mathbf{g}_{\mathcal{P}}\mathbf{g}_{\mathcal{P}}^{\top}]\,\mathbf{e}_{\mathcal{P}}\right].$ (12)
@@
$s=\Delta\ell\approx\mathbb{E}_{\mathbf{x}\sim D}\left[\tfrac{1}{2}\,\mathbf{e}_{\mathcal{P}}^{\top}\,\mathbb{E}[\mathbf{g}_{\mathcal{P}}\mathbf{g}_{\mathcal{P}}^{\top}]\,\mathbf{e}_{\mathcal{P}}\right],$ (13)